OpenStudy (anonymous):

The price of an American Airlines ticket to Boston increased to $600. This is a 15% increase. What was the old fare? (Round to nearest cent.)

OpenStudy (anonymous):

600/1.15

OpenStudy (anonymous):

I'm Trying to help my daughter with her math and haven't done this in a very long time so I take 600 divide it by 1.15

jimthompson5910 (jim_thompson5910):

Let x = old fare New fare is $600, so the difference is 600 - x Divide this by the old fare (x) and set this equal to 0.15 which represents 15% so we have \[\Large \frac{600-x}{x} = 0.15\]

jimthompson5910 (jim_thompson5910):

if you calculate 600/1.15, you get the same answer as you would if you solved \[\Large \frac{600-x}{x} = 0.15\] for x

OpenStudy (anonymous):

So the Awnswer would be 521.73913 rounded to the nearest cent would be 521.74

jimthompson5910 (jim_thompson5910):

correct